Solve by elimination method.
An airplane travels 400 miles against the wind in 5 hours, and makes the return trip with the same wind in 2 hours. Find the rate of the wind.

Answer by Alan3354(69443)


An airplane travels 400 miles against the wind in 5 hours, and makes the return trip with the same wind in 2 hours. Find the rate of the wind.
---------
a = airplane's windspeed
w = windspeed
------------
400/5 = 80 mi/hr upwind
400/2 = 200 mi/hr downwind
------
a + w = 200
a - w = 80
------------------- Add
2a = 280
a = 140
w = 60
=======================
a = average of the 2, = (80 + 200)/2 = 140 mi/hr
w = 200 - 140 = 60 mi/hr